OpenStudy (anonymous):

A cooler contains 5 cans of soda, 2 colas 2 orange and 1 cherry. Two cans are selected at random without replacement. Find the prob that at least one can is orange

OpenStudy (dumbsearch2):

so 5/7 is the answer. Click "Best answer" if it helped. ^_^

OpenStudy (kropot72):

\[P(1)=\frac{\left(\begin{matrix}2 \\ 1\end{matrix}\right)\left(\begin{matrix}8 \\ 1\end{matrix}\right)}{\left(\begin{matrix}10 \\ 2\end{matrix}\right)}=\frac{16}{45}\] \[P(2)=\frac{\left(\begin{matrix}2 \\ 2\end{matrix}\right)\left(\begin{matrix}8 \\ 0\end{matrix}\right)}{\left(\begin{matrix}10 \\ 2\end{matrix}\right)}=\frac{1}{45}\] Probability of at least 1 can is orange is \[\frac{16}{45}+\frac{1}{45}=\frac{17}{45}\]